# When Respondents Don’t Care Anymore: Identifying the Onset of Careless Responding

### v2

@@ Results
RE 0.064 (0.066)
LSP 2.660 (0.753)
(d) Flagged by antonym and personal reliability, but not CODERS (i.e., there is no estimated onset item).
Note. Standard deviations are given in parentheses. The subtables correspond to the four selected participants from Figure 3. Positioning of subtables is consistent with positioning of the subfigures in Figure 3, e.g., Table 2(a) refers to the same participant as Figure 3(a).
First, we focus on Figure 3(a) and Table 2(a). CODERS identifies carelessness onset at item 148, but neither benchmark method flags this respondent as careless. We observe that the average reconstruction error, as well as the variability in the reconstruction errors, is much larger before the estimated onset item than afterward. Since the pre-onset responses could be reconstructed well (except for a brief spike around item 110) while post-onset responses could not, there is evidence that the participant started the questionnaire attentively, but succumbed to survey fatigue about halfway through and subsequently engaged in inconsistent careless respondening.
The second example from Figure 3(b) and Table 2(b) shows another participant that is flagged by CODERS but no benchmark method. The estimated onset of careless responding is at item 58. While the LSP sequence clearly changes around this item, even more excessive values are found for later items, with a maximum LSP value of 31. In fact, only four out of the $n=20,994$ participants in the sample yield more extreme LSP values. Hence, there is strong evidence that this participant responded with (increasingly) lengthy sequences of fixed patterns as the questionnaire progressed. That is, they likely started engaging in invariable careless responding.
The third example, shown in Figure 3(c) and Table 2(c), is similar to the first example in the sense that it suggests an initially attentive respondent starting to give inconsistent careless responses from some point onward in the questionnaire, as indicated by a higher average and a larger variability in the reconstruction errors. Here, the careless onset is estimated at item 119. The crucial difference with the first example is that this participant was not only flagged as careless by CODERS, but also by psychometric antonym. A plausible explanation for this difference may be found in the earlier careless onset. If a participant responds carelessly to a large enough majority of the items, a method designed to detect carelessness throughout the questionnaire (such as the benchmark methods) has a better chance at flagging that participant as careless.
@@
While these four examples are naturally selective, similar plots as in Figure 3 for all participants flagged by CODERS or the two benchmark methods can be found in our online repository at https://github.com/mwelz/carelessonset_plots. In general, two key insights can be derived from the above analysis. First, we find empirical evidence that CODERS can indeed detect various types of partial careless responding, specifically inconsistent and invariable carelessness. Common screeners for careless responding throughout the questionnaire, such as personal reliability or psychometric antonym, are typically only effective for one type of careless responding. Second, CODERS is not a replacement for existing screeners, but rather a complimentary tool that provides new insights if participants display careless responding behavior only in part of the questionnaire.
Comparison with Johnson, (2005)
We now compare our findings with those reported by Johnson, (2005) regarding the incidence of careless responding. This comparison is summarized in Table 3.
Table 3: Incidence of careless responding in the data of Johnson, (2005).
Johnson, (2005) Our findings
Type of careless responding Method Number Method Number
Straightliners Longstring + screeplot 808 (3.6%)
Longstring + cutoff by Costa & McCrae, (2008) 1455 (6.5%)
Inconsistent Reliability 41 (0.2%) Reliability 40 (0.2%)
Antonym 185 (0.8%) Antonym 186 (0.8%)
Either 213 (1.0%) Either 213 (1.0%)
Partial CODERS ($\alpha=0.001$) 158 (0.7%)
CODERS ($\alpha=0.005$) 580 (2.6%)
CODERS ($\alpha=0.010$) 836 (3.7%)
[…]
Johnson, explores two different cutoff values for the longstring index to identify straightliners: one obtained via a screeplot and another suggested by Costa & McCrae, (2008),66endnote: 6Johnson, (2005) refers to a pre-publication version of Costa & McCrae, (2008). with the latter yielding almost twice as many flagged respondents (3.6% vs. 6.5%). As the data set provided by the author already excludes straightliners (among others), we could not attempt to replicate these findings.
Regarding inconsistent careless responding, we find minor differences in the number of flagged respondents compared to Johnson,: 40 vs. 41 for personal reliability, and 186 vs. 185 for psychometric antonym. These minor differences are potentially due to rounding: the psychometric antonym and personal reliability scores provided by Johnson, are rounded by two decimal places, whereas the scores used to compute the reported results in Johnson, (2005) were likely unrounded. As in Johnson, (2005), we find that personal reliability and psychometric antonym agree on 13 of the flagged participants.
Following strategies that “[eliminate] more non-attentive responders, albeit at the cost of more false positives”, Johnson, (2005, p. 116) applies the cutoff value of Costa & McCrae, (2008) for identifying straightliners, and considers respondents that are found by either personal reliability or psychometric antonym as inconsistent careless respondents. This results in a total of 1668 detected careless respondents (7.4%). Employing a similar strategy regarding CODERS, one may choose a significance level of 1%. Adding the corresponding 836 participants yields a total of 2504 flagged careless respondents (11.2%).
Hence, we find evidence that the incidence of careless responding might be substantially underreported in Johnson, (2005) due to hitherto undetected partially careless respondents. Moreover, an incidence of 11.2% in this data set would be higher than the 5–10% that the literature deems problematic for further analysis (e.g., Arias et al.,, 2020; Credé,, 2010).
Simulation Experiments
After demonstrating how CODERS yields new insights in data from a seminal study, we further study the behavior of CODERS via simulation experiments inspired by existing survey measurements.
Data Generation
We generate rating-scale responses of $n=500$ respondents to $p=300$ items. Each item has five Likert-type answer categories (anchored by 1 = “strongly disagree” and 5 = “strongly agree”). The simulated questionnaire measures 30 constructs, each of which is measured by a scale of 10 items. We simulate reliable scales by imposing that items from the same scale have a correlation of 0.7, which results in a Cronbach’s $\alpha$ value of 0.959.77endnote: 7We use package simstudy (Goldfeld & Wujciak-Jens,, 2020) to generate the data, for which the target values of correlations only hold approximately on the population level. Hence, the same applies to Cronbach’s $\alpha$ values. Items from different scales are uncorrelated. All 10 items within a given scale follow the same response probability distribution. We thereby consider the two types of distributions defined in Table 4: centered about the midpoint and skewed towards agreeing. Each of the two distributions is used in 15 scales, yielding the total of $2\times 15\times 10=300$ items. We reverse five items per scale (i.e., half the items) and randomize the order of the items. Note that distributions skewed towards disagreement are implicitly covered by this design because of the item reversal.
Table 4: Response probability distributions of items within different scales.
Type of distribution $P[X=1]$ $P[X=2]$ $P[X=3]$ $P[X=4]$ $P[X=5]$
Centered 0.15 0.20 0.30 0.20 0.15
Skewed 0.10 0.15 0.20 0.25 0.30
Note. The random variable $X$ denotes the response to an item with five Likert-type answer categories.
Subsequently, we select a fixed percentage $\gamma\in\{0,0.02,0.04,0.1,0.2,\dots,0.6,0.8,1\}$ of respondents to be partially careless. That is, $\gamma$ denotes the prevalence of careless responding. For each of the selected partially careless respondents, we sample a carelessness onset item from which onward the remaining responses are careless. We consider two settings for the carelessness onset: (i) early onset, which we randomly sample from the item set $\{30,31,\dots,269,270\}$ so that careless responding starts between 10% and 90% of all items, and (ii) later onset, which we randomly sample from $\{150,151,\dots,269,270\}$, i.e., between half and 90% of all items.
[…]
We repeat the above described data generating process 100 times. Importantly, we keep the location of carelessness onset item of each participant constant across the 100 repetitions, which allows for neat visualization of the results. We compare three variations of CODERS based on (i) both reconstruction errors and the LSP sequence, (ii) reconstruction errors only, and (iii) the LSP sequence only. For the autoencoder, we set the number of nodes in the bottleneck layer equal to the number of measured constructs (i.e., 30).
Performance Evaluation
To quantify how accurately the onset of careless responding is detected, we compute the adjusted Rand index (ARI) (Hubert & Arabie,, 1985, see Appendix D.1 for details). The ARI is best known as a measure of clustering performance: a value of 0 is expected for a random clustering, while a perfectly recovered clustering results in a value of 1. In the careless onset detection problem for a given respondent, we can view the items before the careless onset as one cluster and the items from the onset onwards as another. Specifically, the closer the ARI is to its maximum value 1, the better CODERS performs at accurately locating the true carelessness onset item. On the other hand, an ARI value close to 0 indicates that the estimated onset is far from the true onset. If a respondent is attentive throughout the questionnaire but CODERS incorrectly identifies a changepoint, the ARI takes a value of 0. The same holds if CODERS fails to flag a changepoint for a partially careless respondent.
We report the average ARI across respondents. Note that averaging the ARI over attentive respondents yields the proportion of attentive respondents who are correctly identified as attentive. Conversely, subtracting this number from 1 gives the proportion of attentive respondents that are incorrectly flagged as partially careless.
Results
We start presenting the results of CODERS for the three dimensions by means of a visual inspection. For a carelessness prevalence level of $\gamma=0.6$ and the early onset setting, Figure 4 displays a heatmap holding the indices of the participants in its rows and the item indices in its columns. For participant $i\in\{1,\dots,n\}$ and item $j\in\{1,\dots,p\}$, the color scale of the cell at position $(i,j)$ measures the frequency across the 100 repetitions with which the response of participant $i$ to item $j$ has been flagged as changepoint: a more saturated red indicates that the corresponding item was flagged more frequently as a changepoint. The black rectangles visualize the true location of the careless onset items (recall that those are fixed across the repetitions). The $y$-axis of the plot is rearranged so that the participant indices are ordered according to their respective carelessness onset. Good detection performance is characterized by flagging cells that are close to the corresponding onset items. Indeed, we can see that there is a saturated red band that closely follows the black rectangles. Note that the top 40% of the rows in Figure 4 very rarely see flagged changepoints. Those are the respondents who never become careless. Based on this visual examination, CODERS seems to not only accurately estimate the careless onset, but also discriminate well between partially careless respondents and fully attentive respondents. Results are similar for other combinations of carelessness prevalence and onset, hence we omit the corresponding heatmaps.
Figure 4: Heatmap of carelessness onset detection performance of CODERS.
Note. Of the $n=500$ respondents, 60% are partially careless according to the early onset setting. The significance level of CODERS is set to 0.1%. A more saturated red in a cell corresponds to a higher frequency across the 100 repetitions of the cell being flagged as carelessness onset item ($x$-axis) of a given respondent ($y$-axis). The black rectangles denote the true carelessness onset items (which are constant across the repetitions) and the $y$-axis is rearranged according to the location of true carelessness onset. An absence of a black rectangle (occurs in 40% of all respondents) indicates that a given respondent never becomes careless.
Figure 5: ARI of CODERS for various carelessness prevalence levels.
Note. The ARI is averaged over fully attentive respondents and partially careless respondents (columns) for different carelessness onset settings (rows), and the significance level of CODERS is set to 0.1%.
Varying the prevalence of careless responding, Figure 5 visualizes the ARI averaged over attentive respondents (left column) and partially careless respondents (right column) for the early onset (top row) and later onset (bottom row) settings. Each of the three variations of CODERS rarely flag changepoints in fully attentive respondents, except when carelessness prevalence is extremely high. While the percentage of false positives is close to the significance level 0.1% until 40% prevalence, overrejection remains moderate at 1% for 60% prevalence, before it rises to about 5% for early onset and 10% for later onset at 80% prevalence. For partially careless respondents, the ARI of CODERS with both dimensions remains remarkably constant throughout carelessness prevalence levels at about 0.75 and 0.85 for early and later onset, respectively, mirroring an excellent performance of flagging changepoints in correct locations. In contrast, the ARI using only the LSP sequences amounts to about 0.6 throughout prevalence levels in both onset settings, whereas the ARI using only the reconstruction errors drops from about 0.5 for 2% prevalence to about 0.3 for 100% prevalence with early onset, and from about 0.75 to about 0.35 with later onset. It is important to note that such a drop in performance is expected when using only one indicator of carelessness due to various types of careless responding being simulated. Reconstruction errors are designed to capture inconsistent carelessness and may therefore miss invariant carelessness, while LSP sequences are designed to capture invariant carelessness and may therefore miss inconsistent carelessness.
Figure 6: ARI of CODERS for various carelessness prevalence levels.
Note. The ARI is averaged over the participants corresponding to each careless response type (columns) for different carelessness onset settings (rows), and the significance level of CODERS is set to 0.1%. Note that for straightlining and pattern responding, the results for CODERS using all two dimensions and LongstringPattern only are nearly identical, resulting in overplotted lines.
To explore this further, Figure 6 displays the ARI averaged over subgroups of participants corresponding to each type of careless responding (separate columns) for early onset (top row) and later onset (bottom row). Indeed, reconstruction errors yield much better performance for inconsistent carelessness (random and extreme responding) than for invariable carelessness (straightlining and pattern responding), in particular in the later onset setting. Conversely, using LSP sequences results in near-perfect detection of straightlining and pattern responding with an ARI of about 0.99, but performs particularly poorly for random responding with an ARI close to 0. This highlights the importance of combining the two dimensions in CODERS for accurately detecting various types of partially careless responding.
We now zoom in on the results in Figure 6 for CODERS using reconstruction errors only, which vary by type of (inconsistent) carelessness, as well as by carelessness prevalence and onset location. First, the performance is clearly better for extreme responding than for random responding, with the ARI being in general higher and remaining more stable across prevalence levels. Considering the relatively flat probability distributions of the attentive responses (see Table 4), distinguishing random responses is simply a much more challenging problem. Second, the heterogeneity in performance across prevalence levels and onset location is likely due to the autoencoder requiring sufficiently many attentive responses to successfully distinguish between attentive and (inconsistent) careless responding. If carelessness tends to onset later in the survey, the autoencoder has more information to learn attentive response patterns, making it easier to distinguish between attentive and careless responses. Similar reasoning applies for carelessness prevalence. When fewer responses are careless, the autoencoder has access to more attentive responses from which to learn. In general, these results suggest that CODERS can accurately identify inconsistent careless responding unless careless responses are excessively present in the data through a high prevalence of careless respondents whose carelessness onset occurs early in the survey. However, having an overwhelmingly large number of participants with largely uninformative responses suggests a fundamental issue with data quality, in case of which any subsequent analysis may yield questionable results.
To summarize, this simulation experiment suggests that CODERS succeeds in combining complimentary indicators for carelessness onset in order to capture different manifestations of careless responding.
Additional Simulations
We have conducted a variety of additional simulation experiments, which are described and discussed in Appendix D.2. Specifically, we investigate different choices of the significance level in CODERS, varying correlation structures (and therefore measurement reliability), as well as varying sample sizes. Overall, the findings presented above are representative for those situations. For low to moderate carelessness prevalence (as in the empirical application), choosing a higher significance level (e.g., 1%) leads to an increase in power while retaining size control. On the other hand, for high prevalence levels (about 40% or higher), increasing the significance level yields too many false positives. As expected, power decreases (increases) slightly with decreasing (increasing) measurement reliability or sample size.
We conclude from our simulation experiments that CODERS reliably detects carelessness onset while successfully discriminating between partially careless respondents and fully attentive respondents. In addition, our simulations highlight the importance of using multiple indicators to capture different manifestations of carelessness (cf. Goldammer et al.,, 2020; Huang et al.,, 2012; Meade & Craig,, 2012; Ward & Meade,, 2023).
Extensions, Limitations and Alternatives
Below, we address the potential incorporation of response times into CODERS, and we discuss general limitations as well as potential alternatives.
@@
If per-item response times are available, they can easily be incorporated into CODERS as an additional series to be used in the changepoint detection. If only per-page response times are available and pages are relatively short, an additional series can be obtained by assigning to each response the average page time, i.e., the time the participant spent on the page on which the item is located divided by the number of items on that page (which inspired by the page time index of Bowling et al.,, 2023). If a participant at some point in the survey, for instance, starts responding without reading the items, we expect a changepoint in response time towards considerably faster responses (cf. Bowling et al.,, 2023, 2016; Meade & Craig,, 2012; Huang et al.,, 2012).
@@
Limitations
Throughout this paper, we have assumed that all participants start as truthful and accurate respondents, while some of them may begin to respond carelessly from a certain item onward. However, in certain scenarios, this assumption may not hold true.
First, as we have already seen in our empirical application, some participants may respond carelessly throughout all survey items (cf. Figure 3(d) and Table 2(d)). Consequently, CODERS is unlikely to flag a changepoint due to an absence of a change in behavior. Nevertheless, the literature has established a plethora of methods for detecting respondents who have been careless throughout the questionnaire. In this sense, CODERS—which is intended for detecting partial carelessness—is complementary to those existing methods. If one suspects that respondents who have been careless throughout all items are present, we recommend to apply established detection methods following the guidelines of Arthur et al., (2021) and Ward & Meade, (2023). Respondents that are flagged as careless by such methods, but for whom no changepoint was flagged by CODERS, have likely been careless throughout the entire survey.
Second, there may be some participants who alternate between periods of attentive and careless responding (Berry et al.,, 1992; Clark et al.,, 2003; Meade & Craig,, 2012). In this case, CODERS may only flag one of the possibly multiple changepoints. CODERS should therefore be viewed as a promising first step towards detecting periods of careless responding, with an extension for the detection of multiple changepoints being an area of further research.
Finally, the changepoint detection procedure of Shao & Zhang, (2010) employed by CODERS is based on asymptotic arguments when the number of items is large. Hence, CODERS is designed for lengthy questionnaires, and it is likely that the statistical power drops considerably in short questionnaires. However, partial careless responding due to fatigue may be less prevalent in short questionnaires (cf. Bowling et al.,, 2021). Studying the onset of careless responding in shorter surveys is therefore beyond the scope of this paper.
Potential Alternatives to CODERS
First, the recommendation that one should not include more than three detection items (Meade & Craig,, 2012) is fairly restrictive as carelessness may have onset at a much earlier item than the detection item, resulting in a failure to exclude a potentially large number of careless responses. In principle, a form of carelessness onset detection may be possible by a priori including detection items in the survey, such as bogus, instructed, or self-report items: one may consider a failed detection item (i.e., the participant provides a response that would not be expected from an attentive respondent) as an estimate of the onset of careless responding. However, there are a number of practical and theoretical disadvantages with this approach (see Curran,, 2016, pp. 13–15, for a more general discussion on the caveats of detection items). First, the estimates of the carelessness onset may be crude at best. Recommendations of Meade & Craig, (2012) suggest including at most three bogus, instructed, or self-report items, spaced every 50–100 items. Consequently, the carelessness onset may occur much earlier than at at the detection item, leaving a potentially large number of careless responses undetected. Second, attentive respondents may ironically choose an unexpected response to a bogus item because they find doing so humorous (Meade & Craig,, 2012) or due to misinterpretation of the item (Curran & Hauser,, 2019).88endnote: 8For instance, Curran & Hauser, (2019) find that some respondents would agree to the popular bogus item “All my friends say I would make a great poodle” because they believe that they are loyal friends and they associate loyalty with dogs. Third, a careless respondent may choose the expected response by mere chance or lie in self-report items (DeSimone & Harms,, 2018; Edwards,, 2019), resulting in failure to identify this careless respondent. Fourth, it is unclear if, e.g., the response “somewhat disagree” to a detection item for which “strongly disagree” is expected should be considered careless (Ward & Meade,, 2023). Finally, it may not be possible to include detection items in the survey for legal or ethical reasons, or because the survey analyst has no control over the survey design.
In addition, the software packages careless (Yentes & Wilhelm,, 2021) and ufs (Peters & Gruijters,, 2021) support the calculation of an intra-individual variance (IRV) across subsets of responses.99endnote: 9Specifically, this is implemented in the functions irv() from package careless and carelessReport() from package ufs. If a set of responses towards the end of the questionnaire exhibits a sufficiently larger (Marjanovic et al.,, 2015) or lower (Dunn et al.,, 2018) variance than an earlier set of responses, there might evidence for an onset of partially careless responding. First, these packages only support splitting the responses into subsets of equal length, which does not allow for accurate detection of the carelessness onset. Second, a more general implementation for confirming the presence of partial carelessness in arbitrary subsets would require a priori knowledge on the onset item that defines the subsets, which is unrealistic in practice. Third, it is unclear when a variance can be considered sufficiently low or high to infer carelessness: determining cutoff values below or above which a participant is to be flagged as careless may depend on survey design and survey content.1010endnote: 10For example, if 5-point scales are used a lower cutoff is necessary than for a 9-point scale, since IRVs will be very different for those scales.
Instead of screening for careless responding after data collection, one may attempt to prevent such behavior during survey administration. Preventive measures against careless responding may entail promising rewards for responding accurately (Gibson & Bowling,, 2020), as well as threatening with punishment for careless responding or the presence of a proctor (Bowling et al.,, 2021). There is evidence that each of these three preventive measures can prevent or at least delay carelessness onset (Bowling et al.,, 2021; Gibson & Bowling,, 2020; Huang et al.,, 2012). However, either of these three preventive measures can be undesirable or unethical in practice. Administering proctored questionnaires is often impractical and expensive, and is typically infeasible in online studies. Arthur et al., (2021) and Bowling et al., (2021) advise against using threats if one wishes to maintain a positive long-term relationship with questionnaire participants, such as when there is a professional relationship between researcher and participants. Moreover, threats seem inappropriate in longitudinal studies or when the participants are members of a questionnaire panel (as is common in survey research), or when they are suspected to suffer from anxiety, depression, or trauma (as is common, e.g., in research on wellbeing in the workplace, or more general in psychological research). On the other hand, promising rewards may cross ethical boundaries in organizational surveys and in the context of employment-related decisions (Arthur et al.,, 2021).
Notwithstanding, we stress that preventive methods against carelessness are complementary to detection methods. If preventative methods are appropriate, combining them with screening techniques such as CODERS may further enhance quality and reliability of the collected data for subsequent analysis.

### v3

@@ Results
RE 0.079 (0.105)
LSP 2.813 (1.115)
(f) Presumably inconsistent throughout the survey (no changepoint item detected by CODERS).
Note. Standard deviations are given in parentheses. The subtables correspond to the selected participants from Figure 11. Positioning of subtables is consistent with positioning of the subfigures in Figure 11, e.g., Table 3(a) refers to the same participant as Figure 11(a).
First, we focus on the two participants in Figures 11(a)11(b) and Tables 3(a)3(b). CODERS identifies carelessness onset at item 148 and item 246, respectively. We observe that the average reconstruction error, as well as the variability in the reconstruction errors, is much larger before the estimated onset item than afterward. Since the pre-onset responses could (generally) be reconstructed well while post-onset responses could not, there is evidence that these participants succumbed to survey fatigue and eventually engaged in inconsistent careless responding.
These two cases are representative for the majority of respondents flagged by CODERS, likely starting attentively but starting to respond carelessly at some point. In addition, more respondents flagged by CODERS seem to exhibit inconsistency than invariability, as Johnson, already removed straightliners from the publicly available data set. However, CODERS also identifies participants with remarkably different behavior.
The participant in Figure 11(c) and Table 3(c) shows similar behavior as the first two examples, albeit in reverse. CODERS identifies a changepoint at item 63, with reconstruction errors being large in terms of magnitude and variability before the changepoint while being small afterward. We conclude that this participant may have started the survey as an inconsistent careless respondent, but then shifted toward attentive responding.
For the participant in Figure 11(d) and Table 3(d), CODERS identifies a changepoint at item 58. Both the reconstruction errors and the LSP sequence clearly change around this item, with even more excessive LSP values occurring in later items (maximum value of 31). In fact, only four out of the $n=20,994$ participants in the sample yield more extreme LSP values. Furthermore, the reconstruction errors before the changepoint are of comparable magnitude and variability to those of the previous examples of inconsistent responding. Hence, there is strong evidence that this participant initially responded inconsistently, but switched to invariable responding characterized by (increasingly) lengthy sequences of fixed patterns as the questionnaire progressed.
Next, we discuss the example in Figure 11(e) and Table 3(e), for whom CODERS detects a changepoint at item 142. It turns out that this respondent repeated the pattern 5-5-3-3-3 in the first half of the survey. Even after the changepoint, the LSP values are among the highest in the sample, with the participant switching between the similar patterns 1-1-3-3-3, 5-1-3-3-3, and 1-5-3-3-3 for the remainder of the questionnaire. Although there is a clear change in behavior, as found by CODERS, this participant seems to respond invariably throughout the survey.
@@
In summary, we reiterate that the first two examples are representative for the majority of participants flagged by CODERS, in line with the assumption that respondents start attentive and possibly become careless. Although other cases indicate that this assumption is too restrictive in practice, CODERS—which processes item-level indicators of inconsistency and invariability for changepoints—successfully uncovered relevant changes in response behavior. In light of these findings, we proceed to discuss limitations and potential extensions.
Limitations and Extensions
We start by reviewing general limitations of CODERS together with (possible) remedies, before discussing the potential incorporation of response times into CODERS.
Limitations
First, some participants may respond carelessly throughout all survey items (cf. Figure 11(f) and Table 3(f)). Consequently, CODERS is unlikely to flag a changepoint due to an absence of a change in behavior. Nevertheless, the literature has established a plethora of methods for detecting careless respondents. In this sense, CODERS—which is intended for detecting partial carelessness through a change in behavior—is complementary to those existing methods. If one suspects that respondents who have been careless throughout all items are present, we recommend to apply established detection methods following the guidelines of Arthur et al., (2021) and Ward & Meade, (2023). Respondents that are flagged as careless by such methods, but for whom no changepoint was flagged by CODERS, have likely been careless throughout the survey.
Second, a detected changepoint does not necessarily indicate a change from attentive responding to careless responding. A participant may respond carelessly in the beginning of the survey and switch to responding attentively later on (cf. Figure 11(c) and Table 3(c)), or they may switch between careless response behaviors through the course of the survey (cf. Figures 11(d)11(e) and Tables 3(d)3(e)). For such cases, our empirical application demonstrates that CODERS—combined with a post-hoc analysis of the identified segments of responses—remains a valuable tool for gaining detailed insights into careless response behaviors.
Third, there may be participants who alternate between periods of attentive and careless responding (Berry et al.,, 1992; Clark et al.,, 2003; Meade & Craig,, 2012). Our simulations indicate that CODERS remains conservative in this case and rarely flags one of the changepoints (see Appendix C.5). CODERS should therefore be viewed as a promising first step towards detecting periods of careless responding, with an extension for the detection of multiple changepoints being an area of further research.
Finally, the changepoint detection procedure of Shao & Zhang, (2010) employed by CODERS is based on asymptotic arguments when the number of items is large. Hence, the number of items in the questionnaire is the main determinant for the statistical power of CODERS. For the detection of inconsistent careless responding, the number of participants plays a secondary role regarding power: the more attentive responses the autoencoder can learn from, the better the reconstruction errors can distinguish between attentive and careless responding (cf. Appendix C.3). The LSP sequence of a given participant, on the other hand, depends only on their own responses, meaning that the number of observations should not play a role concerning power in detecting invariable careless responding. CODERS is thus primarily designed for relatively lengthy questionnaires, and our simulations suggest that there is a gradual drop in statistical power for detecting inconsistency as survey length decreases (see Appendix C.2). However, partial carelessness due to survey fatigue may be less prevalent in short questionnaires (cf. Bowling et al.,, 2021). A detailed study of partially careless responding in shorter surveys is therefore beyond the scope of this paper.
@@
If per-item response times are available, they can easily be incorporated into CODERS as an additional series to be used in the changepoint detection. If only per-page response times are available and pages are relatively short, an additional series can be obtained by assigning to each response the average page time, i.e., the time the participant spent on the page on which the item is located divided by the number of items on that page (which is inspired by the page time index of Bowling et al.,, 2023). If a participant at some point in the survey, for instance, starts responding without reading the items, we expect a changepoint in response time towards considerably faster responses (cf. Bowling et al.,, 2023, 2016; Meade & Craig,, 2012; Huang et al.,, 2012).